OpenStudy (anonymous):

Use Newton's method to find xsubscript6, when xsubscript0 = -7, for the function f(x) = x^2 + 13.5x + 45.

OpenStudy (anonymous):

Place this in Physics section....

OpenStudy (amistre64):

in other words, start at x=-7 and work thru the tangent line equations 6 times ....

OpenStudy (amistre64):

all this is meant for is practice to get you into the feel of the Newton method

OpenStudy (anonymous):

\[x _{i+1} = x_{i} - f(x_{i})/f'(x_{i})\] Since \[x_{0} = -7\] and f'(x) = 2x + 13.5 \[x_{0+1} = -7 - f(-7)/f'(-7) = -7 - ( (-7)^{2} + 13.5(-7) + 45)/(2(-7) + 13.5)\] That means that \[x_{1} = -7 - (49-94.5+45)/(-14+13.5)\] So \[x_{1} = -8\] To find \[x_{2}\] remember that \[x_{i+1} = x_{i} - f(x)/f'(x)\] so \[x_{2} = x_{1} - f(x_{1})/f'(x_{1})\] where \[x_{1} = -8\] Continue to this process until you find \[x_{6}\]